# Grokfast: Accelerated Grokking by Amplifying Slow Gradients

### v1

@@ 1 Introduction
Grokking is a recently discovered phenomenon where generalization is achieved long after a model overfits to the training data. The phenomenon was first reported by Power et al. [2022] for a two-layer Transformer [Vaswani et al., 2017] trained using a simple algorithmic dataset. Later, Liu et al. [2022b] have shown that similar artifacts are observed for various model architectures trained with a variety of datasets, including images, languages, and graphs. Many theory-oriented works have tried to justify the effect by relating the grokking phenomenon to the previously known double descent phenomenon [Humayun et al., 2023, Davies et al., 2023], yet its cause and sufficient conditions have not been fully characterized.
@@ 2.4 Limitations
Although Algorithm 1 shows highly effective results, it requires $w$ times more memory to store all the previous gradients, limiting its utilization. Replication of the model parameters also makes the training slower; using $w=100\,$, the training time per iteration is increased by ×$\times 2.4\,$ measured with a single 2080 Ti GPU. Still, the reduction of wall clock time before the delayed generalization of the results in Figure 7 is ×$\times 21.0\,$, which is also a notable reduction of time. Though the computation time does not linearly scale with the memory requirements, Algorithm 1 is not generally applicable to the larger models. This problem leads to our new design utilizing sequential filter in the next section.
@@ 4.1 Algorithmic Data
We also conduct ablation studies to find out the effect of hyperparameters $\lambda\,$, $\alpha\,$, and weight decay for our Grokfast algorithm with an EMA filter. The optimal hyperparameters are found with grid search as in Section 2. Figures LABEL:fig:ema:acc_gain through LABEL:fig:ema:acc_wd summarizes the results. Recalling that our main idea is at the design of a low-pass filter, the momentum parameter $\alpha$ of Algorithm 2 as well as the window size parameter $w$ of Algorithm 1 are equivalent to the cutoff frequency of the underlying filter. Experiments in Figures LABEL:fig:ema:acc_gain through LABEL:fig:ema:acc_wd as well as those in Figures 4 and 7 show that there exists a sweet spot in cutoff frequency that corresponds to the generalization-inducing gradient signal. From our empirical studies, we recommend $\lambda\in[0.5,5]$ and $\alpha\in[0.8,0.99]\,$. The weight decay is, like in typical optimization problems, dependent on the task of interest.
@@ 4.2 MNIST
Besides the simple algorithmic reasoning task, where the data is relatively simple, Liu et al. [2022b] report the similar delayed generalization can also be observed in many conventional tasks if the model goes through a special treatment. To demonstrate the generalizability of our Grokfast modification of the optimization process, we try to accelerate the speed of generalization under those reported models and tasks. The first is a three-layer ReLU-MLP trained for MNIST classification task [Deng, 2012] which exhibits the grokking phenomenon. Figure 10 summarizes the results, showing that our Algorithm 2 successfully accelerate the delayed generalization. With $\alpha=0.85\,$, $\lambda=0.2\,$, and $\mathrm{wd}=20\,$, the delay until grokking is reduced by ×$\times 22.0\,$. Moreover, the final evaluation accuracy becomes higher from 89.5% to 91.5%.
@@ 6 Related Work
The recently discovered grokking phenomenon [Power et al., 2022] signifies one possibility of overparameterized neural networks generalizing (and reasoning) beyond memorization of the given dataset. Most of the works thereafter, focuses on identifying its mechanism. Humayun et al. [2023] associated grokking with the double descent phenomenon, wherein the validation error initially increases and then decreases with the expansion of model parameters [Nakkiran et al., 2021, Belkin et al., 2019]. Their assertion posits that the model captures slow, generalizing patterns, underscoring the critical role of proper optimization. Interestingly, while weight decay amplifies the double descent effect [Pezeshki et al., 2022], it contributes to enhanced generalization in grokking scenarios [Power et al., 2022]. Liu et al. [2022b] found more examples of grokking in various tasks and analyzed their training mechanism through loss landscape analysis. Thilak et al. [2022] found similarity between grokking and the slingshot mechanism in adaptive optimizers. Barak et al. [2022] argued that optimizers reach delayed generalization by amplifying sparse solutions through hidden progress. Regularizers such as weight decay [Nanda et al., 2023] and the choice of the optimizer [Liu et al., 2022a] are highlighted as an important factor of training a model that groks. Our work is a continuation of this discussion by providing a generalizable tool for practical study of grokking phenomenon. Through our discussion, we suggest state transition model of the grokking, and visualize the trajectory of the model weights in the parameter space during training.

### v2

@@ 1 Introduction
Grokking is a recently discovered phenomenon where generalization is achieved long after a model overfits to the training data. The phenomenon was first reported by Power et al. [2022] for a two-layer Transformer [Vaswani et al., 2017] trained using a simple algorithmic dataset. Later, Liu et al. [2022b] have shown that similar artifacts are observed for various model architectures trained with a variety of datasets, including images, languages, and graphs. Many theory-oriented works have tried to justify the effect by relating the grokking phenomenon to the previously known double descent phenomenon [Davies et al., 2023, Huang et al., 2024], yet its cause and sufficient conditions have not been fully characterized.
@@ 2.4 Limitations
Although Algorithm 1 shows highly effective results, it requires $w$ times more memory to store all the previous gradients, limiting its utilization. Replication of the model parameters also makes the training slower; using $w=100\,$, the training time per iteration is increased by ×$\times 2.4\,$ measured with a single 1080 Ti GPU. Still, the reduction of wall clock time before the delayed generalization of the results in Figure 7 is ×$\times 20.5\,$, which is also a notable reduction of time. Though the computation time does not linearly scale with the memory requirements, Algorithm 1 is not generally applicable to the larger models. This problem leads to our new design utilizing sequential filter in the next section.
@@ 4.1 Algorithmic Data
We also conduct ablation studies to find out the effect of hyperparameters $\lambda\,$, $\alpha\,$, and weight decay for our Grokfast algorithm with an EMA filter. The optimal hyperparameters are found with grid search as in Section 2. Figures LABEL:fig:ema:acc_gain through LABEL:fig:ema:acc_wd summarizes the results. Recalling that our main idea is at the design of a low-pass filter, the momentum parameter $\alpha$ of Algorithm 2 as well as the window size parameter $w$ of Algorithm 1 are equivalent to the cutoff frequency of the underlying filter. Experiments in Figures LABEL:fig:ema:acc_gain through LABEL:fig:ema:acc_wd as well as those in Figures 4 and 7 show that there exists a sweet spot in cutoff frequency that corresponds to the generalization-inducing gradient signal. From our empirical studies, we recommend $\lambda\in[0.1,5]$ and $\alpha\in[0.8,0.99]\,$. The weight decay is, like in typical optimization problems, dependent on the task of interest.
@@ 4.2 MNIST
Besides the simple algorithmic reasoning task, where the data is relatively simple, Liu et al. [2022b] report the similar delayed generalization can also be observed in many conventional tasks if the model goes through a special treatment. To demonstrate the generalizability of our Grokfast modification of the optimization process, we try to accelerate the speed of generalization under those reported models and tasks. The first is a three-layer ReLU-MLP trained for MNIST classification task [Deng, 2012] which exhibits the grokking phenomenon. Figure 10 summarizes the results, showing that our Algorithm 2 successfully accelerate the delayed generalization. With $\alpha=0.8\,$, $\lambda=0.1\,$, and $\mathrm{wd}=2.0\,$, the delay until grokking is reduced by ×$\times 22.0\,$. Moreover, the final evaluation accuracy becomes higher from 89.8% to 91.2%.
@@ 6 Related Work
The recently discovered grokking phenomenon [Power et al., 2022] signifies one possibility of overparameterized neural networks generalizing (and reasoning) beyond memorization of the given dataset. Most of the works, thereafter, focus on identifying its mechanism. Recent studies associated grokking with a double descent phenomenon in the DNN mapping geometry training dynamics [Humayun et al., 2023], the speed of pattern learning  [Davies et al., 2023], and the sizes of models and datasets [Huang et al., 2024], wherein the validation error initially increases and then decreases with the expansion of model parameters [Nakkiran et al., 2021, Belkin et al., 2019]. To investigate the internal roles of each model component during grokking, Nanda et al. [2023] employed mechanistic interpretability, a technique from the XAI domain, and revealed that grokking may not occur abruptly but rather exhibits an internal progress measure. Their assertion posits that the model captures slow, generalizing patterns, underscoring the critical role of proper optimization. Interestingly, while weight decay amplifies the double descent effect [Pezeshki et al., 2022], it contributes to enhanced generalization in grokking scenarios [Power et al., 2022]. Liu et al. [2022b] found more examples of grokking in various tasks and analyzed their training mechanism through loss landscape analysis. Thilak et al. [2022] found a similarity between grokking and the slingshot mechanism in adaptive optimizers. Barak et al. [2022] argued that optimizers reach delayed generalization by amplifying sparse solutions through hidden progress. Regularizers such as weight decay [Nanda et al., 2023] and the choice of the optimizer [Liu et al., 2022a] are highlighted as important factors in training a model that groks. Our work is a continuation of this discussion by providing a generalizable tool for the practical study of the grokking phenomenon. Through our discussion, we suggest a state transition model of the grokking and visualize the trajectory of the model weights in the parameter space during training.
@@ B.4 IMDb
Appendix C Autograd Implementation
We have argued that our implementation of Algorithm 1 and 2 costs only a few additional lines of code. We demonstrate this by presenting the exact code we developed with the PyTorch [Paszke et al., 2019] autograd package. The readers who are interested can also consult our official implementation [^2] https://github.com/ironjr/grokfast.
Algorithms 1 and 2 are implemented as follows:
1# Grokfast-MA (Algorithm 1)
2def gradfilter_ma(
3 m: nn.Module,
4 grads: Optional[Dict[str, deque]] = None,
5 window_size: int = 100,
6 lamb: float = 5.0,
7 filter_type: Literal[’mean’, ’sum’] = ’mean’,
8 warmup: bool = True,
9) -> Dict[str, deque]:
10 if grads is None:
11 grads = {n: deque(maxlen=window_size) for n, p in m.named_parameters() if p.requires_grad}
12
13 for n, p in m.named_parameters():
14 if p.requires_grad:
15 grads[n].append(p.grad.data.detach())
16
17 if not warmup or len(grads[n]) == window_size:
18 if filter_type == "mean":
19 avg = sum(grads[n]) / len(grads[n])
20 elif filter_type == "sum":
21 avg = sum(grads[n])
22 else:
23 raise ValueError(f"Unrecognized filter_type {filter_type}")
24 p.grad.data = p.grad.data + avg * lamb
25
26 return grads
27
28# Grokfast (Algorithm 2)
29def gradfilter_ema(
30 m: nn.Module,
31 grads: Optional[Dict[str, torch.Tensor]] = None,
32 alpha: float = 0.98,
33 lamb: float = 2.0,
34) -> Dict[str, torch.Tensor]:
35 if grads is None:
36 grads = {n: p.grad.data.detach() for n, p in m.named_parameters() if p.requires_grad}
37
38 for n, p in m.named_parameters():
39 if p.requires_grad:
40 grads[n] = grads[n] * alpha + p.grad.data.detach() * (1 - alpha)
41 p.grad.data = p.grad.data + grads[n] * lamb
42
43 return grads
This helper method can be applied to any optimization framework involving the autograd package by inserting a single line between the calculation of the gradients and the optimizer call as follows:
1# ... any initialization code before starting the training loop.
2grads = None
3
4# Training loop.
5for batch in dataloader:
6 model.zero_grad()
7 output = model(batch)
8 loss = criteria(output)
9
10 # Calculate the gradients.
11 loss.backward()
12
13 # Option 1: Grokfast (has argument alpha, lamb)
14 grads = gradfilter_ema(model, grads=grads, alpha=alpha, lamb=lamb)
15 # Option 2: Grokfast-MA (has argument window_size, lamb)
16 # grads = gradfilter_ma(model, grads=grads, window_size=window_size, lamb=lamb)
17
18 # Call the optimizer.
19 optimizer.step()
20
21 # ... any additional logging codes.
Note that line 2 and line 14 in the code above are the only modification we made.
Appendix D Time and Memory Requirements
Table 3: Quantitative results of Grokfast with a Transformer decoder trained for the algorithmic data (modular multiplication). The experiments corresponds to that of Figure 2 and LABEL:fig:ema:acc & LABEL:fig:ema:loss.
Algorithm Iterations @ 95% Val. Acc. Wall Clock Time @ 95% Val. Acc. (s) VRAM (MB) Latency Per Iteration (s)
Baseline $39890$ $5984$ $290$ $0.15$
[…]
Table 4: Quantitative results of Grokfast with an MLP trained for MNIST (Figure 10).
Algorithm Iterations @ 95% Val. Acc. Wall Clock Time @ 95% Val. Acc. (s) VRAM (MB) Latency Per Iteration (ms)
Baseline $44022$ $1928$ $196$ $43.8$
[…]
Table 5: Quantitative results of Grokfast with a G-CNN trained for QM9 (Figure 10).
Algorithm Minimum Val. Loss VRAM (MB) Latency Per Iteration (ms)
Baseline $0.00659$ $216$ $40.2$
Grokfast $0.00348$ $216$ $41.4$
Table 6: Quantitative results of Grokfast with an LSTM trained for IMDb (Figure 11).
Algorithm Best Val. Acc. Minimum Val. Loss VRAM (MB) Latency Per Iteration (ms)
Baseline $0.84$ $0.517$ $754$ $20.4$
Grokfast $0.90$ $0.412$ $762$ $21.2$
This section delivers additional demonstration of the efficiency of our Grokfast algorithm. As we have argued in Section 2.4 and Section 3, the additional computational burden from our augmentation is compensated by the larger-scale acceleration of the delayed generalization. The additional cost of VRAM memory is also negligible compared the baseline. The time and the memory requirements in the Tables 6 through 6 are measured with a single GTX 1080 Ti GPU.
Appendix E More Visualization
Figure 13: Normalized trajectories of model parameters from five runs of the experiment of Figure 2. The baseline optimization algorithm without Grokfast guide the model to the overfitting states (state B) relatively closer to the initialization states (state A). After reaching the overfitting state, however, the model parameters travel far away to reach the generalization states (state C). Grokfast instead guide the model parameters to the alternative generalization states (state C), which are much closer to the initialization states (state A).
[…]
We first notice that the overfitting states (state B) from each of the two optimizers are clearly different. The baseline algorithm without Grokfast reaches the overfitting states (state B), which are relatively nearer to the initialization states (state A) than those of Grokfast algorithm. However, as soon as the model overfits, the weights continue to deviate far from the points where overfitting first occured (state B). As a result, the final generalization (state C) happens much far away from the initialized weights (state A). It is notable that the five generalization states (state C) from different instances of the baseline optimizer vary significantly. The difference between the states C of the baseline is much larger than that of the baseline’s overfitting states (state B) and that of the states B and C from our Grokfast algorithm. In contrast, the training dynamics from Grokfast results in a distinct set of trajectories that lead to the generalization states (state C) much closer to the initial weights. Moreover, difference within the trajectories from Grokfast is much smaller than that of the baseline algorithms. This conclusion is also verifiable from Figure LABEL:fig:traj:mag and Table 2 in a quantitative manner.